The perimeter of a regular hexagon is 96 cm find the length of each side
Answer by jim_thompson5910(35256)

A regular hexagon has 6 equal sides. At this point, each side is unknown. We'll call each side x for now

We'll use the formula Perimeter = 6*(side length) and the given perimeter to find the side length

Perimeter = 6*(side length)
P = 6*x
96 = 6x
6x = 96
6x/6 = 96/6 ... divide both sides by 6 to isolate x
x = 16


So each side is 16 cm long.